The CFO of the supermarket chain Aldi is examining an investment in a new delivery service for internet orders. The initial investment of the project consists of a cash outflow of €700,000 for the following items: a) distribution fleet, b) computer servers, c) delivery packages, and d) other fixed assets. These assets will be fully depreciated for tax purposes over a 5-year period, which is the lifetime of the project, using the straight-line depreciation method. At the end of this project, the distribution fleet and computer servers can be sold for €60,000.

The revenue of the new service will be equal, for the first year, to 15% of the current total sales of the traditional service, which are €7.5 million. For the following years, the revenues of the new service will be 20%, 25%, 30%, and 35% respectively of the initial traditional sales. This new service for internet ordering is a premium service. Therefore, the prices of the home delivery goods will be 5% higher than the current supermarket prices.

It should be taken into account that the cost of goods sold represents 70% of the supermarket sales price. Aldi expects to spend each year 7% of the sales of the new service for advertising. The company already has the available warehouse capacity needed for this service. The service will occupy 10% of the warehouse. The warehouse is rented at an annual cost of €300,000. The wages for the project workers will be €150,000 per year, however, 20% of them are workers transferred from other Aldi businesses (the company has a lifetime job policy). The distribution costs (fuel, etc.) are 3% of sales of the new service.

The introduction of the new internet ordering service will have as a consequence the reduction of the traditional sales of the supermarket business. Sales in the traditional supermarket business are expected to decrease from €7.5 million a year to €7.4 million a year for the next five years.

Accounts payable are equal to 4% of cost of goods sold, inventory corresponds to 8% of cost of goods sold, and accounts receivables are 10% of sales. At the end of the internet sales project, all remaining net working capital would be liquidated, and the cash recovered. Aldi has profits from its current business, pays taxes of 25% and has a cost of capital equal to 10%

. You are required to:

a. Calculate the project’s profit after tax throughout its five-year cycle

b. What is the change in Net Working Capital for each year?

c. Calculate the project’s Net Cash Flows for each year

d. Evaluate the project according to its NPV, IRR, Profitability Index, and Payback Period

Ι NEED EXPLANATION FOR EVERY STEP..THANK YOU!
Answer #1
Present Value(PV) of Cash Flow:
(Cash Flow)/((1+i)^N)
i=discount rate=Cost of Capital=10%=0.10
N=Year   of Cash Flow
Loss of Contribution from existing Business €                 30,000 (7.5-7.4) Million *(1-0.7)
After Tax Salvage Value=60000*(1-0.25) €                 45,000
C=(A-B)/10 Annual depreciation €              140,000 (700000/5)
Year 1 2 3 4 5
D Revenue as percentage of current Sales 15% 20% 25% 30% 35%
E=D*7.5 million Expected Revenue €              1,125,000 €           1,500,000 €           1,875,000 €           2,250,000 €           2,625,000
N Year 0 1 2 3 4 5
a Initial investmentCash Flow €               (700,000)
b Annual Revenue €           1,125,000 €           1,500,000 €           1,875,000 €           2,250,000 €           2,625,000
c=b*70% Cost of goods sold €            (787,500) €        (1,050,000) €        (1,312,500) €        (1,575,000) €        (1,837,500)
d Depreciation Expenses €            (140,000) €            (140,000) €            (140,000) €            (140,000) €            (140,000)
e Cost of Warehouse space(10%*300000) €              (30,000) €              (30,000) €              (30,000) €              (30,000) €              (30,000)
15% 1,125,000 € 25% 1,875,000 20% ,500,000 € 21 30% 2,250,000 € 35% ,625,000 1 € 2 17 N 1 1,125,000 € (787,500) € (140,000) €G44 X fi =IRR(G38:138) H € (140,000) € (30,000) € (120,000) € (33,750) € (30,000) € (78,750) € (95,000 € (71,250) € 140,000 €